OpenStudy (anonymous):

x and y intercepts of -2y=3x-6

OpenStudy (saifoo.khan):

When x intercepts y = 0 When y intercepts x = 0

OpenStudy (saifoo.khan):

getting it?

OpenStudy (anonymous):

no.

OpenStudy (saifoo.khan):

When x intercepts y = 0 -2y=3x-6 -2(0) = 3x - 6 0 = 3x - 6 3x = 6 x = 6/3 x = 2 (2,0) When y intercepts x = 0 -2y=3x-6 -2y = 3(0) - 6 -2y = -6 2y = 6 y = 6/2 y = 3 (0,3)

OpenStudy (anonymous):

i get it thanks